A HIGH RESOLUTION FAR INFRARED STUDY OF LOW LYING VIBRATIONAL BANDS OF 3-OXETANONE

\underline{ZIQIU CHEN} AND JENNIFER VAN WIJNGAARDEN, Department of Chemistry, University of Manitoba, Winnipeg MB R3T 2N2 Canada.

Rotationally resolved vibrational spectra of the four lowest frequency bands of the four-membered heterocycle 3-oxetanone (c-C_{3}H_{4}O_{2}) have been collected with a resolution of 0.00096 cm^{−1} using the far infrared beamline at the Canadian Light Source synchrotron. The modes observed correspond to motions best described as: ring puckering v_{21} at 139.55 cm^{−1}, C=O deformation (out-of-plane) v_{20} at 399.59 cm^{−1}, C=O deformation (in-plane) v_{16} at 448.19 cm^{−1}, and the ring deformation v_{7} at 684.99 cm^{−1}. The observed spectra exhibit rich spectroscopic features including hot bands originating from the low energy ring puckering band and a first order a-type Coriolis interaction between the closely spaced v_{16} and v_{20} bands. The assignment and analysis of these dense spectra will be presented.